# Packing subgraphs in regular graphs

### v1

@@ 5.1 Finding vertex-disjoint super-regular subgraphs that cover all vertices outside $V_{0}$
By (G’1)(G’3) and Lemma 3.3, the graph $G^{\prime}$ has a Hamilton cycle, which we denote by $C=(v_{1}\ldots v_{n^{\prime}})$, where $n^{\prime}=|V(G^{\prime})|$. Let $w:E(\Gamma)\to\mathbb{R}^{\geq 0}$ be the edge-weighting where $w(V_{i}V_{j})$ is the number of edges in $C$ with one end in $V_{i}$ and the other in $V_{j}$, divided by $2|V_{1}|$. We claim that $w$ is a perfect fractional matching in $\Gamma$; that is, $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$. Indeed, for every $i\in[m]$, since $G^{\prime}$ has no edges with both ends in $V_{i}$, the number of edges in $C$ with an end in $V_{i}$ is exactly $2|V_{i}|=2|V_{1}|$, so $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$, as claimed. By Fact˜5.4, this implies that $\Gamma$ has a perfect $2$-matching, as required. ∎
@@
By Proposition˜3.7, we can move $3\varepsilon|U_{i}|$ vertices from $U_{i}$ (for every $i\in[2m]$) into the set $V_{0}$, so that the resulting subgraphs $G^{\prime}[U_{2i-1},U_{2i}]$ are $(4\varepsilon,\mu/4)$-super-regular for every $i\in[m]$. By moving less than $2t$ additional vertices from each $U_{i}$ to $V_{0}$, we also assume that $|U_{i}|$ is divisible by $2t$ for every $i\in[2m]$. We claim that after moving these vertices, the following properties hold.
@@ 5.2 Splitting the super-regular subgraphs
For every $i\in[2m]$, let $\left\{U_{i}^{(1)},\ldots,U_{i}^{(5)}\right\}$ be a random partition of $U_{i}$ into sets of sizes $\xi|U_{i}|$, $\xi|U_{i}|$, $(1/2t)|U_{i}|$, $(2/3)|U_{i}|$, $(1/3-1/2t-2\xi)|U_{i}|$, respectively. Then (U1)(U4) hold by definition. Moreover, by Lemma˜3.1 and a union bound, (U5) holds with probability at least 1o(1)1-o(1); fix an outcome such that it holds.
@@ 5.3 Covering $V_{0}$ using a well-distributed collection of copies of $K_{t,t}$
Fix a subset $W\subseteq N(v,U^{\prime})$ of size $(\xi\zeta/16)n$ and write $Z\coloneqq U^{\prime}-W$. Since every vertex $u$ in $G$ satisfies $d_{G}(u,U^{\prime})\geq(\xi\zeta/8)n$, every vertex in $W$ has at least $(\xi\zeta/16)n$ neighbours in $Z$. By Lemma˜3.8, the graph $G[W,Z]$ has a $K_{t,t}$, say, with vertex set $A\cup B$ where $A\subseteq W$ and $B\subseteq Z$. Let $B^{\prime}$ be a subset of $B$ of size t1t-1. Then $G[A,B^{\prime}\cup\{v\}]$ is a copy of $K_{t,t}$ in $G[V_{0}\cup U^{(1)}]$, which is vertex-disjoint from $V(\mathcal{K}_{1})$ and contains exactly one vertex from $V_{0}$, contradicting the maximality of $\mathcal{K}_{1}$. ∎
Next, we find a small number of vertex-disjoint $K_{t,t}$’s in $G[U^{(2)}]$ that will allow us to achieve the desired divisibility condition. Recall that the support of a function $f:A\to B$, denoted $\operatorname{supp}(f)$, is the set of elements $a$ in $A$ with $f(a)\neq 0$.
@@
Let $W$ be a walk in $\Gamma_{\mathrm{ref}}$ as in the previous paragraph, and write $W=U_{i_{0}},\ldots,U_{i_{2\ell}}$, where $U_{i_{0}}=U_{a}$ and $U_{i_{2\ell}}=U_{b}$. Since $W$ visits each vertex in $\Gamma_{\mathrm{ref}}$ at most three times and $|U_{i}\cap V(\mathcal{K}^{\prime})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|\leq 6tm$ for $i\in[2m]$, we can apply Lemma˜3.8 to iteratively choose pairwise vertex-disjoint $K_{t,t}$’s in $U^{(2)}-V(\mathcal{K}^{\prime})$, denoted $K_{1},\ldots,K_{\ell}$, such that for every $j\in[\ell]$, $K_{j}$ consists of $t$ vertices in $U^{(2)}_{i_{2j-1}}$, $f(a)$ vertices in $U^{(2)}_{i_{2j-2}}$ and tf(a)t-f(a) vertices in $U^{(2)}_{i_{2j}}$. Write $\mathcal{K}_{2}\coloneqq\mathcal{K}^{\prime}\cup\{K_{1},\ldots,K_{\ell}\}$. It is easy to check that $\mathcal{K}_{2}$ satisfies the requirements of the induction hypothesis since $|U_{i}\cap V(\mathcal{K}_{2})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|+3t\leq 3t|\operatorname{supp}(f)|$ and $|U_{i}\cap V(\mathcal{K}_{2})|\equiv f(i)\!\pmod{t}$ for $i\in[2m]$. This completes the induction step and proves the claim. ∎
Let $\mathcal{K}_{1}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim˜5.6. For $i\in[2m]$, let $f(i)\coloneqq-|U_{i}\cap V(\mathcal{K}_{1})|\pmod{t}=|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$. Then
@@
The first equality follows by the definition of $f$, and the second equality follows from the fact that $Y-V_{0}=\bigcup_{i\in[m]}U_{2i}$. For the third equality, first recall that $|U_{1}|=\ldots=|U_{2m}|$ by (G’4), so $|X-V_{0}|=|Y-V_{0}|$. This implies that $|X\cap V_{0}|=|Y\cap V_{0}|$ since $G$ is bipartite and balanced. Combining this with the fact that $V_{0}\subseteq V(\mathcal{K}_{1})$, and that every copy of $K_{t,t}$ in $\mathcal{K}_{1}$ contains exactly one vertex of $V_{0}$, we get $|(X-V_{0})\cap V(\mathcal{K}_{1})|=|(Y-V_{0})\cap V(\mathcal{K}_{1})|$ from which the third equality follows easily. The final equality follows from $|V_{0}|$ and $|V(\mathcal{K}_{1})|$ being divisible by $2t$. Using (17), (18) and Claim˜5.7, we can find a collection $\mathcal{K}_{2}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G[U^{(2)}]$ satisfying $|U_{i}^{(2)}\cap V(\mathcal{K}_{2})|\equiv f(i)\equiv-|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$ for $i\in[2m]$, which implies that $|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})|$ is divisible by $t$, as desired.
@@ 5.4 Covering the remaining vertices of $G$ with vertex-disjoint copies of $K_{t,t}$
It follows that $H$ has average degree at least $d^{\prime}(1+\eta)-32\eta t\cdot|V(H)|$, as required for (b). Notice that (c) is the same as (U6) and (d) is the same as (U7). For proving (e), note that $|V(H)-V(H^{\prime})|\leq|U^{(1)}|+|U^{(2)}|\leq 2\xi n\leq 8t\xi\cdot|V(H)|$. Finally, for proving (f), suppose that $G$ is bipartite and balanced, and denote its bipartition by $\{X,Y\}$. Recall that, by (G’4), $|V_{0}\cap X|=|V_{0}\cap Y|$ (since $|X|=|Y|$). By the choice of $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ (see Claim˜5.6 and Claim˜5.7), we have $|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap X)|=|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap Y)|$, and thus
@@
Thus, by applying Lemma˜3.3 with $H$, $d^{\prime}(1+\eta)$, $c/2t$, $32\eta t$, $8t\xi$, $\gamma/(256t^{2})$, $\zeta/20$, $V(H^{\prime}), V(H)V(H)$ playing the roles of $G$, $d$, $c$, $\eta$, $\xi$, $\gamma$, $\zeta$, $W$, $n$ respectively, we find that there is a Hamilton cycle in $H^{\prime}$, which contains a perfect matching in $H^{\prime}$ (since $|V(H^{\prime})|$ is even), thus proving the claim. ∎
Let $\mathcal{M}^{\prime}$ be a perfect matching in $H^{\prime}$ (guaranteed by Claim˜5.8). For $i,j\in[2m]$, denote by $f(i,j)$ the number of edges in $\mathcal{M}^{\prime}[U_{i}^{\prime(3)},U_{j}^{\prime(3)}]$.
@@
Now recall that, by (G’6), for all $i,j\in[2m]$, the graph $G^{\prime}[U_{i},U_{j}]$ is $5\varepsilon$-regular, and it has density more than $\mu/5$ if it has at least one edge. Hence, we have $e_{G^{\prime}}(U_{i}^{\prime\prime},U_{j}^{\prime\prime})>(\mu/5-5\varepsilon)\cdot|U_{i}^{\prime\prime}|\cdot|U_{j}^{\prime\prime}|$, so by the Kővári–Sós–Turán theorem (Lemma˜3.8), there is a $K_{t,t}$ in $G^{\prime}[U_{i}^{\prime\prime},U_{j}^{\prime\prime}]$, contradicting the maximality of $\mathcal{K}_{3}$ and proving the claim. ∎
Let $\mathcal{K}_{3}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim˜5.9. For $i\in[2m]$, write $U_{i}^{\prime}\coloneqq U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})$. For $i\in[2m]$, by the choice of $U_{i}^{\prime(3)}$ as in (19) and the fact that $|U_{i}\cap V(\mathcal{K}_{3})|=t\cdot\sum_{j\in[2m]}f(i,j)=t\cdot|U_{i}^{\prime(3)}|$, we have
@@
Recall that the sets $U_{i}$ for $i\in[2m]$ have the same size, which is divisible by $2t$ by (G’4). Hence, the sets $U^{\prime}_{i}$ for $i\in[2m]$ have the same size, which is divisible by $t$. Moreover, by (U4), (U5), (G’5) and the fact that $U_{i}^{(5)}$ is a subset of $U_{i}^{\prime}$ for each $i\in[2m]$, we have that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ is $(10\varepsilon,\mu/50)$-super-regular for $i\in[m]$. Therefore, it follows from the blow-up lemma (Lemma˜3.6) that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ has a perfect $K_{t,t}$-packing (since $\varepsilon\ll\mu,1/t$). These perfect $K_{t,t}$-packings of $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ for $i\in[m]$, along with $\mathcal{K}_{1}$, $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$, form a perfect $K_{t,t}$-packing in $G$. This completes the proof of Lemma 5.1. ∎
6 $H$-packings in dense regular graphs
In this section, we combine Lemma˜4.1 and Corollary˜5.2 to prove Theorem 1.4, which asserts that for every bipartite graph $H$ and every constant $0<c\leq 1$, there is a constant $C>0$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$, has an $H$-packing that covers all but at most $C$ vertices of $G$. Denote the sizes of the two parts of $H$ by $a$ and $b$, and write $t\coloneqq a+b$. Then, it is easy to see that there are two vertex-disjoint copies of $H$ covering all the vertices of the complete bipartite graph $K_{t,t}$. Hence, Theorem 1.4 follows immediately from the following.

### v2

@@ 5.1 Finding vertex-disjoint super-regular subgraphs that cover all vertices outside $V_{0}$
By (G’1)(G’3) and Lemma 3.3, the graph $G^{\prime}$ has a Hamilton cycle, which we denote by $C=(v_{1}\ldots v_{n^{\prime}})$, where $n^{\prime}=|V(G^{\prime})|$. Let $w:E(\Gamma)\to\mathbb{R}^{\geq 0}$ be the edge-weighting where $w(V_{i}V_{j})$ is the number of edges in $C$ with one end in $V_{i}$ and the other in $V_{j}$, divided by $2|V_{1}|$. We claim that $w$ is a perfect fractional matching in $\Gamma$; that is, $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$. Indeed, for every $i\in[m]$, since $G^{\prime}$ has no edges with both ends in $V_{i}$, the number of edges in $C$ with an end in $V_{i}$ is exactly $2|V_{i}|=2|V_{1}|$, so $\sum_{j\in[m]}w(V_{i}V_{j})=1$ for every $i\in[m]$, as claimed. By Fact 5.4, this implies that $\Gamma$ has a perfect $2$-matching, as required. ∎
@@
By Proposition 3.7, we can move $3\varepsilon|U_{i}|$ vertices from $U_{i}$ (for every $i\in[2m]$) into the set $V_{0}$, so that the resulting subgraphs $G^{\prime}[U_{2i-1},U_{2i}]$ are $(4\varepsilon,\mu/4)$-super-regular for every $i\in[m]$. By moving less than $2t$ additional vertices from each $U_{i}$ to $V_{0}$, we also assume that $|U_{i}|$ is divisible by $2t$ for every $i\in[2m]$. We claim that after moving these vertices, the following properties hold.
@@ 5.2 Splitting the super-regular subgraphs
For every $i\in[2m]$, let $\left\{U_{i}^{(1)},\ldots,U_{i}^{(5)}\right\}$ be a random partition of $U_{i}$ into sets of sizes $\xi|U_{i}|$, $\xi|U_{i}|$, $(1/2t)|U_{i}|$, $(2/3)|U_{i}|$, $(1/3-1/2t-2\xi)|U_{i}|$, respectively. Then (U1)(U4) hold by definition. Moreover, by Lemma 3.1 and a union bound, (U5) holds with probability at least 1o(1)1-o(1); fix an outcome such that it holds.
@@ 5.3 Covering $V_{0}$ using a well-distributed collection of copies of $K_{t,t}$
Fix a subset $W\subseteq N(v,U^{\prime})$ of size $(\xi\zeta/16)n$ and write $Z\coloneqq U^{\prime}-W$. Since every vertex $u$ in $G$ satisfies $d_{G}(u,U^{\prime})\geq(\xi\zeta/8)n$, every vertex in $W$ has at least $(\xi\zeta/16)n$ neighbours in $Z$. By Lemma 3.8, the graph $G[W,Z]$ has a $K_{t,t}$, say, with vertex set $A\cup B$ where $A\subseteq W$ and $B\subseteq Z$. Let $B^{\prime}$ be a subset of $B$ of size t1t-1. Then $G[A,B^{\prime}\cup\{v\}]$ is a copy of $K_{t,t}$ in $G[V_{0}\cup U^{(1)}]$, which is vertex-disjoint from $V(\mathcal{K}_{1})$ and contains exactly one vertex from $V_{0}$, contradicting the maximality of $\mathcal{K}_{1}$. ∎
5.4 Constructing copies of $K_{t,t}$ to ensure divisibility
Next, we find a small number of vertex-disjoint $K_{t,t}$’s in $G[U^{(2)}]$ that will allow us to ensure that the number of uncovered vertices in every $U_{i}$ is divisible by $t$. Recall that the support of a function $f:A\to B$, denoted $\operatorname{supp}(f)$, is the set of elements $a$ in $A$ with $f(a)\neq 0$.
@@ 5.4 Constructing copies of $K_{t,t}$ to ensure divisibility
Let $W$ be a walk in $\Gamma_{\mathrm{ref}}$ as in the previous paragraph, and write $W=U_{i_{0}},\ldots,U_{i_{2\ell}}$, where $U_{i_{0}}=U_{a}$ and $U_{i_{2\ell}}=U_{b}$. Since $W$ visits each vertex in $\Gamma_{\mathrm{ref}}$ at most three times and $|U_{i}\cap V(\mathcal{K}^{\prime})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|\leq 6tm$ for $i\in[2m]$, we can apply Lemma 3.8 to iteratively choose pairwise vertex-disjoint $K_{t,t}$’s in $U^{(2)}-V(\mathcal{K}^{\prime})$, denoted $K_{1},\ldots,K_{\ell}$, such that for every $j\in[\ell]$, $K_{j}$ consists of $t$ vertices in $U^{(2)}_{i_{2j-1}}$, $f(a)$ vertices in $U^{(2)}_{i_{2j-2}}$ and tf(a)t-f(a) vertices in $U^{(2)}_{i_{2j}}$. Write $\mathcal{K}_{2}\coloneqq\mathcal{K}^{\prime}\cup\{K_{1},\ldots,K_{\ell}\}$. It is easy to check that $\mathcal{K}_{2}$ satisfies the requirements of the induction hypothesis since $|U_{i}\cap V(\mathcal{K}_{2})|\leq 3t\cdot|\operatorname{supp}(f^{\prime})|+3t\leq 3t|\operatorname{supp}(f)|$ and $|U_{i}\cap V(\mathcal{K}_{2})|\equiv f(i)\!\pmod{t}$ for $i\in[2m]$. This completes the induction step and proves the claim. ∎
Let $\mathcal{K}_{1}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim 5.6. For $i\in[2m]$, let $f(i)\coloneqq-|U_{i}\cap V(\mathcal{K}_{1})|\pmod{t}=|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$. Then
@@
The first equality follows by the definition of $f$, and the second equality follows from the fact that $Y-V_{0}=\bigcup_{i\in[m]}U_{2i}$. For the third equality, first recall that $|U_{1}|=\ldots=|U_{2m}|$ by (G’4), so $|X-V_{0}|=|Y-V_{0}|$. This implies that $|X\cap V_{0}|=|Y\cap V_{0}|$ since $G$ is bipartite and balanced. Combining this with the fact that $V_{0}\subseteq V(\mathcal{K}_{1})$, and that every copy of $K_{t,t}$ in $\mathcal{K}_{1}$ contains exactly one vertex of $V_{0}$, we get $|(X-V_{0})\cap V(\mathcal{K}_{1})|=|(Y-V_{0})\cap V(\mathcal{K}_{1})|$ from which the third equality follows easily. The final equality follows from $|V_{0}|$ and $|V(\mathcal{K}_{1})|$ being divisible by $2t$. Using (18), (19) and Claim 5.7, we can find a collection $\mathcal{K}_{2}$ of pairwise vertex-disjoint copies of $K_{t,t}$ in $G[U^{(2)}]$ satisfying $|U_{i}^{(2)}\cap V(\mathcal{K}_{2})|\equiv f(i)\equiv-|U_{i}^{(1)}\cap V(\mathcal{K}_{1})|\pmod{t}$ for $i\in[2m]$, which implies that $|U_{i}\cap V(\mathcal{K}_{1}\cup\mathcal{K}_{2})|$ is divisible by $t$, as desired.
@@ 5.5 Covering the remaining vertices of $G$ with vertex-disjoint copies of $K_{t,t}$
It follows that $H$ has average degree at least $d^{\prime}(1+\eta)-32\eta t\cdot|V(H)|$, as required for (b). Notice that (c) is the same as (U6) and (d) is the same as (U7). For proving (e), note that $|V(H)-V(H^{\prime})|\leq|U^{(1)}|+|U^{(2)}|\leq 2\xi n\leq 8t\xi\cdot|V(H)|$. Finally, for proving (f), suppose that $G$ is bipartite and balanced, and denote its bipartition by $\{X,Y\}$. Recall that, by (G’4), $|V_{0}\cap X|=|V_{0}\cap Y|$ (since $|X|=|Y|$). By the choice of $\mathcal{K}_{1}$ and $\mathcal{K}_{2}$ (see Claim 5.6 and Claim 5.7), we have $|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap X)|=|V(G^{\prime})\cap(V(\mathcal{K}_{1}\cup\mathcal{K}_{2})\cap Y)|$, and thus
@@
Thus, by applying Lemma 3.3 with $H$, $d^{\prime}(1+\eta)$, $c/2t$, $32\eta t$, $8t\xi$, $\gamma/(256t^{2})$, $\zeta/20$, $V(H^{\prime}), V(H)V(H)$ playing the roles of $G$, $d$, $c$, $\eta$, $\xi$, $\gamma$, $\zeta$, $W$, $n$ respectively, we find that there is a Hamilton cycle in $H^{\prime}$, which contains a perfect matching in $H^{\prime}$ (since $|V(H^{\prime})|$ is even), thus proving the claim. ∎ Let $\mathcal{M}^{\prime}$ be a perfect matching in $H^{\prime}$ (guaranteed by Claim 5.8). For $i,j\in[2m]$, denote by $f(i,j)$ the number of edges in $\mathcal{M}^{\prime}[U_{i}^{\prime(3)},U_{j}^{\prime(3)}]$.
@@
Now recall that, by (G’6), for all $i,j\in[2m]$, the graph $G^{\prime}[U_{i},U_{j}]$ is $5\varepsilon$-regular, and it has density more than $\mu/5$ if it has at least one edge. Hence, we have $e_{G^{\prime}}(U_{i}^{\prime\prime},U_{j}^{\prime\prime})>(\mu/5-5\varepsilon)\cdot|U_{i}^{\prime\prime}|\cdot|U_{j}^{\prime\prime}|$, so by the Kővári–Sós–Turán theorem (Lemma 3.8), there is a $K_{t,t}$ in $G^{\prime}[U_{i}^{\prime\prime},U_{j}^{\prime\prime}]$, contradicting the maximality of $\mathcal{K}_{3}$ and proving the claim. ∎
Let $\mathcal{K}_{3}$ be a collection of copies of $K_{t,t}$ as guaranteed by Claim 5.9. For $i\in[2m]$, write $U_{i}^{\prime}\coloneqq U_{i}-V(\mathcal{K}_{1}\cup\mathcal{K}_{2}\cup\mathcal{K}_{3})$. For $i\in[2m]$, by the choice of $U_{i}^{\prime(3)}$ as in (20) and the fact that $|U_{i}\cap V(\mathcal{K}_{3})|=t\cdot\sum_{j\in[2m]}f(i,j)=t\cdot|U_{i}^{\prime(3)}|$, we have
@@
Recall that the sets $U_{i}$ for $i\in[2m]$ have the same size, which is divisible by $2t$ by (G’4). Hence, the sets $U^{\prime}_{i}$ for $i\in[2m]$ have the same size, which is divisible by $t$. Moreover, by (U4), (U5), (G’5) and the fact that $U_{i}^{(5)}$ is a subset of $U_{i}^{\prime}$ for each $i\in[2m]$, we have that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ is $(10\varepsilon,\mu/50)$-super-regular for $i\in[m]$. Therefore, it follows from the blow-up lemma (Lemma 3.6) that $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ has a perfect $K_{t,t}$-packing (since $\varepsilon\ll\mu,1/t$). These perfect $K_{t,t}$-packings of $G^{\prime}[U_{2i-1}^{\prime},U_{2i}^{\prime}]$ for $i\in[m]$, along with $\mathcal{K}_{1}$, $\mathcal{K}_{2}$ and $\mathcal{K}_{3}$, form a perfect $K_{t,t}$-packing in $G$. This completes the proof of Lemma 5.1. ∎
6 $F$-packings in dense regular graphs
In this section, we combine Lemma 4.1 and Corollary 5.2 to prove Theorem 1.4, which asserts that for every bipartite graph $F$ and every constant $0<c\leq 1$, there is a constant $C>0$ such that every $d$-regular graph $G$ of order $n$, with $d\geq cn$, has an $F$-packing that covers all but at most $C$ vertices of $G$. Denote the sizes of the two parts of $F$ by $a$ and $b$, and write $t\coloneqq a+b$. Then, it is easy to see that there are two vertex-disjoint copies of $F$ covering all the vertices of the complete bipartite graph $K_{t,t}$. Hence, Theorem 1.4 follows immediately from the following.